1. Using u=XY solve the boundary valued problem

du/dx + 2du/dy = 4u

u(x,0) = 2e^3x + 4e^5x
1
Expert's answer
2020-04-15T08:38:26-0400

Given :

"\partial u/\partial x+2\partial u/\partial y=4u" ---(1)

Now, let us assume that "u(x,y)=X(x)Y(y)"

Thus; "\partial u/\partial x=X'Y; \partial u/\partial y=XY'" ---(2)

Using (2) in (1);

"\implies X'Y+2XY'=4XY"

"\implies (X'/4X)+(Y'/2Y)=1"

"\implies (X'/4X)=1-(Y'/2Y)"

As LHS is a function of x alone and RHS is a function of y alone.

Then, LHS=RHS=constant=k(say)

"\implies (X'/4X)=1-(Y'/2Y)=k"


Solving "(X'\/4X)=k" ; we get;

"\int X'dx/X=\int 4kdx \implies lnX=4kx+c" ; where c is constant of integration.

"\implies X(x)=Ae^{4kx}" ; ---(3)

A is an arbitrary constant.

Again solving "1-(Y'\/2Y)=k" ; we get;

"\int Y'dy/Y=\int 2(1-k)dy \implies lnY=2(1-k)y+c" ; where c is constant of integration.

"\implies Y(y)=Be^{2(1-k)y}" ---(4)

B is an arbitrary constant.


Using (3) and (4), we get;

"u(x,y)=X(x)Y(y)=(Ae^{4kx})(Be^{2(1-k)y})"

"\implies u(x,y)=Ce^{4kx+2(1-k)y}" ; ---(5)

where C=AB is a arbitrary constant.

The boundary condition is incorrect, there cannot be two terms.

If "u(x,0)=2e^{3x}" ; then

(5) "\implies 2e^{3x}=Ce^{4kx} \implies C=2 ; k=3/4"

Thus, the final solution is ;

"\implies u(x,y)=2e^{3x+y/2}" (Answer)